# Structural and rigidity properties of Lie skew braces

### v1

@@ 2.3. Post-Lie algebras
    $[\zeta,\eta]_{\circ}-[\zeta,\eta]_{\cdot}=\zeta\triangleright\eta-\eta\triangleright\zeta$,
@@
    $\zeta\triangleright[\eta,\gamma]_{\cdot}=[\zeta\triangleright\eta,\gamma]_{\cdot}-[\eta,\zeta\triangleright\gamma]_{\cdot}$,
@@
    $[\zeta,\eta]_{\circ}\triangleright\gamma=\zeta\triangleright(\eta\triangleright\gamma)-\eta\triangleright(\zeta\triangleright\gamma)$.
@@
Notice that by (1) in Definition 2.7 the Lie algebra structure $[\cdot,\cdot]_{\circ}$ is uniquely determined by the pair $([\cdot,\cdot]_{\cdot},\triangleright)$. Hence, one can also define (see, e.g. [2] and [6]) a PLA as a triple $(\mathfrak{g},[\cdot,\cdot]_{\cdot},\triangleright)$, where $(\mathfrak{g},[\cdot,\cdot]_{\cdot})$ is a real Lie algebra and
@@
    $\zeta\triangleright[\eta,\gamma]_{\cdot}=[\zeta\triangleright\eta,\gamma]_{\cdot}-[\eta,\zeta\triangleright\gamma]_{\cdot}$,
@@
$[\zeta,\eta]_{\circ}:=[\zeta,\eta]_{\cdot}+\zeta\triangleright\eta-\eta\triangleright\zeta.$
Indeed, is not hard to verify that the quadruple $(\mathfrak{g},[\cdot,\cdot]_{\cdot},[\cdot,\cdot]_{\circ},\triangleright)$ is indeed a PLA.
An important class of examples of post-Lie algebras is given by pre-Lie algebras (also known as a left-symmetric algebra), which arise when the Lie bracket $[\cdot,\cdot]_{\cdot}$ is trivial. In this setting, the post-Lie product defines a compatible Lie algebra structure via antisymmetrization.
@@
Given a LSB $(G,\cdot,\circ)$, one can associate two Lie algebras:
@@
A post-Lie algebra $(\mathfrak{g},[\cdot,\cdot]_{\cdot},[\cdot,\cdot]_{\circ},\triangleright)$ is integrable if there exists a LSB $(G,\cdot,\circ)$ such that
@@
Consider the pre-Lie algebra $\left(\mathbf{M}_{n}(\mathbb{R}),[\cdot,\cdot]_{\cdot}=0,[\cdot,\cdot]_{\circ},\ \triangleright\right)$, where $[X,Y]_{\circ}:=XY-YX$ is the usual matrix commutator, and the post-Lie product is defined by
@@
$[x,y]_{\cdot}=0,\qquad[x,y]_{\circ}=xy-yx=0,$
so the underlying Lie algebras are abelian. The post-Lie product becomes $x\triangleright y=xy$, the standard multiplication in $\mathbb{R}$. Suppose, for contradiction, that this structure integrates to a LSB which is necessarily $(\mathbb{R},+,+)$. For such a trivial brace, we have
@@ 2.4. Affine actions
    There exists a Lie group structure $\cdot_{G}$ on $G$ such that $(G,\cdot_{G},\circ_{G})$ is a LSB;
@@
(1) $\Rightarrow$ (2): Suppose $(G,\cdot_{G},\circ_{G})$ is a LSB. Then the map
@@
where $\rho(g)=(x,f)$ and $e_{K}$ denotes the identity element of $(K,\cdot_{K})$. It follows that $\rho(G)$ is a regular subgroup of $\mathrm{Aff}(K,\cdot_{K})$. Now, by Proposition 2.4, every regular Lie subgroup of $\mathrm{Aff}(K,\cdot_{K})$ corresponds to a LSB structure $(K,\cdot_{K},\circ_{K})$, and the group $\rho(G)$ is isomorphic to $(K,\circ_{K})$. Let $\Psi:\rho(G)\to(K,\circ_{K})$ be such an isomorphism. Then the composition
@@
which turns $G$ into a Lie group and makes $(G,\cdot_{G},\circ_{G})$ a LSB isomorphic to $(K,\cdot_{K},\circ_{K})$. The result follows. ∎
@@
    For every connected and simply-connected non-linear Lie group $(K,\cdot_{K})$, there exists a real connected linear $(G,\circ)$ and a simply transitive affine action $\rho:(G,\circ_{G})\longrightarrow\mathrm{Aff}(K,\cdot_{K})$.
@@
where $\mathrm{Aut}_{0}(G,\cdot)$ denotes the connected component of the identity in the group of Lie group automorphisms of $G$. Since $G$ is simply-connected, by [23, Thm. 3.18.13] it admits a Levi-Malcev decomposition $(G,\cdot)=R\rtimes S,$ where $R$ is the radical of $G$ and $S$ is a semisimple subgroup. Thus, in the terminology in [33], the subgroup $R$ is a nucleus, i.e. a closed connected and simply-connected normal, solvable Lie subgroup of $G$ such that $G/R$ is reductive. Since $R$ is invariant under the action of $\mathrm{Aut}_{0}(G)$, it follows by [33, Lemma 2] that there exists a faithful representation of $G$ which can be extended to a (possibly non-faithful) representation of the semidirect product $(G,\cdot)\rtimes\mathrm{Aut}_{0}(G,\cdot)$. Now consider the short exact sequence
$1\rightarrow(G,\cdot)\rightarrow(G,\cdot)\rtimes\mathrm{Aut}_{0}(G)\rightarrow\mathrm{Aut}_{0}(G)\rightarrow 1$
Since $(G,\cdot)$ admits a faithful linear representation that extends to the semidirect product, and $\mathrm{Aut}_{0}(G,\cdot)$ is linear, being a Lie subgroup of $\mathrm{GL}(\mathfrak{g})$, where $\mathfrak{g}=\mathrm{Lie}(G,\cdot)$, we may apply [33, Lemma 1] to conclude that $\mathrm{Aff}_{0}(G,\cdot)$ is linear. ∎
Our second lemma shows how a LSB structure can be lifted to the universal covering of a connected LSB.
@@
Let $(G,\cdot,\circ)$ be a connected LSB with identity element $e\in G$. Let $\widetilde{G}$ be the universal covering space of $G$, and let $\pi:\widetilde{G}\rightarrow G$ denote the covering map. Then, for any $\tilde{e}\in\pi^{-1}(e)$, there exists a LSB structure $(\widetilde{G},\tilde{\cdot},\tilde{\circ})$ with identity $\tilde{e}$ such that $\pi$ is a LSB homomorphism from $(\widetilde{G},\tilde{\cdot},\tilde{\circ})$ to $(G,\cdot,\circ)$.
@@
By standard results in Lie group theory (see, e.g., [32, Section 3.24]), there exist two Lie group structures on $\widetilde{G}$, denoted by $(\widetilde{G},\cdot)$ and $(\widetilde{G},\circ)$, both having $\tilde{e}$ as identity element, such that $\pi$ is a Lie group homomorphism from $(\widetilde{G},\cdot)$ to $(G,\cdot)$ and from $(\widetilde{G},\circ)$ to $(G,\circ)$ (here with a slight abuse of notation we are denoting the operations on $\widetilde{G}$ with the same symbol $\cdot$ and $\circ$ used for $G$).
To show that $(\tilde{G},\cdot,\circ)$ is a LSB and that $\pi$ is a LSB homomorphism, we must verify that the identity (1) is satisfied, namely
$[\tilde{a}\circ(\tilde{b}\cdot\tilde{c})]\cdot[(\tilde{a}\circ\tilde{b})\cdot\tilde{a}^{-1}\cdot(\tilde{a}\circ\tilde{c})]^{-1}=\tilde{e}$ (7)
@@
i.e. $\pi\tilde{f}=\pi c_{\tilde{e}}=e$. Indeed, for $c_{\tilde{e}}$ is obvious while for $\tilde{f}$:
@@
with $a=\pi(\tilde{a}),b=\pi(\tilde{b})$ and $c=\pi(\tilde{c})$, where the second equality follows by the fact that $\pi$ is an homomorphism from $(\widetilde{G},\cdot)$ to $(G,\cdot)$ and from $(\widetilde{G},\circ)$ to $(G,\circ)$ and the third equality follows by the fact that $(G,\cdot,\circ)$ is a LSB and so (1) holds true. Moreover, $\tilde{f}$ and $c_{\tilde{e}}$ agree at the point $(\tilde{e},\tilde{e},\tilde{e})$, i.e. $c_{\tilde{e}}(\tilde{e},\tilde{,}e,\tilde{e})=\tilde{f}(\tilde{e},\tilde{e},\tilde{e})=\tilde{e}$ (the second equality follows by sobstituting $(\tilde{e},\tilde{e},\tilde{e})$ into (7)). By uniqueness of lifts on covering spaces (since the space involved are connected), it follows that $\tilde{f}=c_{\tilde{e}}$, and we are done. ∎
@@
Thanks to the correspondence between LSBs and simply transitive affine actions established in Theorem 2.12, Lemma 3.2 admits the following geometric counterpart: Let $\rho:(G,\circ)\;\longrightarrow\;\mathrm{Aff}(K,\cdot)$ be a simply transitive affine action of a connected Lie group $(G,\circ)$ on another connected Lie group $(K,\cdot)$. Then the universal coverings $\widetilde{G}$ and $\widetilde{K}$ carry a lifted action $\widetilde{\rho}:(\widetilde{G},\circ)\;\longrightarrow\;\mathrm{Aff}(\widetilde{K},\cdot),$ which is again simply transitive.
@@
(S1) Let $(G,\cdot,\circ)$ be a LSB, with $(G,\cdot)$ connected simply-connected and linear. By Proposition 2.4, there exists a regular subgroup $H\leq\mathrm{Aff}(G,\cdot)$ such that $H\cong(G,\circ)$. Since $G$ is connected, it follows that $H\leq\mathrm{Aff}_{0}(G,\cdot)$. As $\mathrm{Aff}_{0}(G,\cdot)$ is linear (by Lemma 3.1), $H$ is also linear. Therefore, $(G,\circ)$ is linear as well.
@@
(R2) Let $(G,\cdot,\circ)$ be a connected Lie skew brace such that $(G,\circ)$ is semisimple and let $(\mathfrak{g},[\cdot,\cdot]_{\cdot},[\cdot,\cdot]_{\circ})$ be the corresponding PLA. Then the Lie algebra $(\mathfrak{g},[\cdot,\cdot]_{\circ})$ is semisimple as well. By [12, Th. 3.3], it follows that $(\mathfrak{g},[\cdot,\cdot]_{\cdot})$ is isomorphic to $(\mathfrak{g},[\cdot,\cdot]_{\circ})$ as Lie algebras. Since $(G,\cdot)$ and $(G,\circ)$ are diffeomorphic as manifolds, we conclude that $(G,\cdot)\cong(G,\circ)$ as Lie groups. ∎
@@
and define $G:=H/N$, denoting the group operation on $G$ by $\circ$. We represent cosets in $G$ as $[x,y,\theta]$, with $\theta\in\mathbb{S}^{1}$. As shown in [23, Section 4.8], the group $G$ is not linear. Now consider the following subgroups of $G$:
@@
It is readily verified that $G_{1}$ is a closed normal subgroup of $G$, $G_{2}$ is a closed subgroup, and $G_{1}\cap G_{2}=\{e\}$. Therefore, $G$ is the semidirect product $G=G_{1}\rtimes G_{2}$. Now endow the same underlying set $G=(\mathbb{R}\times\mathbb{S}^{1})\times\mathbb{R}$ with the direct product operation $\cdot$. According to Example 4, the pair of operations (,)($\cdot$,$\circ$)( ⋅ , ∘ ) defines a connected but non-simply-connected LSB, where the group $(G,\cdot)$ is linear, although $(G,\circ)$ is not. This shows that the assumption of simple-connectedness is crucial for the validity of point (S1).
@@
Because $G_{0}$ is connected, condition (S2) of Theorem 1.1 tells us that the group $(G_{0},\circ)$ is solvable. Next, $(G,\cdot)$ is compact and nilpotent, and $G_{0}$ is an open subgroup, so $G_{0}$ has finite index in $(G,\cdot)$. Therefore $|\overline{G}|<\infty$, and $(\overline{G},\cdot,\circ)$ is a finite left skew brace. Nilpotency is preserved by quotients, hence $(\overline{G},\cdot)$ is finite and nilpotent. By [29, Cor. 2.23] the multiplicative group $(\overline{G},\circ)$ is solvable. Since the class of solvable groups is closed under extensions, the group $(G,\circ)$ is also solvable. This completes the proof  [^2] Motivated by the discrete left skew brace setting (cf. [24]), one may call a subset $I\subseteq(G,\cdot,\circ)$ an ideal if it is a closed Lie subgroup satisfying (i) $I\trianglelefteq(G,\cdot)$, (ii) $I\trianglelefteq(G,\circ)$, (iii) $\lambda_{a}(I)=I$ for every $a\in G$. The argument above shows that $G_{0}$ is indeed such an ideal. The closedness of $I$ is precisely what guarantees that the quotient $G/G_{0}$ carries a natural manifold structure.. ∎
@@
(F1) Let $(G,\cdot)$ be a connected, simply connected, non-linear Lie group. Our objective is to endow $G$ with a second Lie group structure $\circ$ such that $(G,\cdot,\circ)$ forms a LSB, and the group $(G,\circ)$ is linear.
@@
In order to get (9) we use again the Levi–Malcev decomposition. So let $R\trianglelefteq G$ be a connected and simply connected solvable Lie subgroup, and $S\leq G$ be a connected and simply connected semisimple Lie, such that
@@
Since $KAN=S\leq G$, we have $ANK=KAN$, so $r\in KAN$. But $R\cap KAN=\{e\}$, thus $r=e$, implying $uv=an\in AN$. However, $uv\in K$, and $K\cap AN=\{e\}$, hence $uv=e$, and so $x=e$. Therefore, $G_{1}\cap G_{2}=\{e\}$. Next, we prove that both $G_{1}$ and $G_{2}$ are linear. Note that $G_{1}=RUV=(RU)V.$ Since $R\trianglelefteq G$ and $R\cap U=\{e\}$, we have $RU\cong R\rtimes U$, which is connected, simply connected, and solvable. Also, $V$ is compact and therefore real linearly reductive. Since $UV$ normalizes $R$ it follows that, it follows that $RU\trianglelefteq G_{1}$, and hence $G_{1}=(RU)\rtimes V$. By [21, Th. 16.2.7], such a group is linear. Moreover, $G_{2}=AN$ is connected, simply connected, and solvable, so it is linear by the same result.
(F2) Let $(G,\circ)$ be a connected and simply connected solvable Lie group. Combining Theorem 2.12 with Theorem 2.11, one can construct a LSB $(G,\cdot,\circ)$ such that the Lie group $(G,\cdot)$ is nilpotent.
@@
To justify each entry of the table in Theorem 1.1, we examine every pair $(G,\cdot)$ and $(G,\circ)$ in turn, constructing explicit non-trivial LSB structures for the boxes marked ✓ and proving non-existence or triviality for those marked with a dash (--).

### v3

@@ 2.3. Post-Lie algebras
    $\zeta,\eta]_{\circ}-[\zeta,\eta]_{\cdot}=\zeta\triangleright\eta-\eta\triangleright\zeta$,
@@
    $\zeta\triangleright[\eta,\gamma]_{\cdot}=[\zeta\triangleright\eta,\gamma]_{\cdot}{+}[\eta,\zeta\triangleright\gamma]_{\cdot}$,
@@
    $\zeta,\eta]_{\circ}\triangleright\gamma=\zeta\triangleright(\eta\triangleright\gamma)-\eta\triangleright(\zeta\triangleright\gamma)$.
@@
Notice that by (1) in Definition 2.8 the Lie algebra structure $\cdot$$\cdot$$_{\circ}$ is uniquely determined by the pair $([\cdot,\cdot]_{\cdot},\triangleright)$. Hence, one can also define (see, e.g. [2] and [6]) a PLA as a triple $(\mathfrak{g},[\cdot,\cdot]_{\cdot},\triangleright)$, where $(\mathfrak{g},[\cdot,\cdot]_{\cdot})$ is a real Lie algebra and
@@
    $\zeta\triangleright[\eta,\gamma]_{\cdot}=[\zeta\triangleright\eta,\gamma]_{\cdot}{+}[\eta,\zeta\triangleright\gamma]_{\cdot}$,
@@
$\zeta,\eta]_{\circ}:=[\zeta,\eta]_{\cdot}+\zeta\triangleright\eta-\eta\triangleright\zeta.$
Indeed, it is not hard to verify that the quadruple $(\mathfrak{g},[\cdot,\cdot]_{\cdot},[\cdot,\cdot]_{\circ},\triangleright)$ is a PLA.
An important class of examples of post-Lie algebras is given by pre-Lie algebras (also known as left-symmetric algebras), which arise when the Lie bracket $\cdot$$\cdot$$_{\cdot}$ is trivial. In this setting, the post-Lie product defines a compatible Lie algebra structure via antisymmetrization.
@@
Given an LSB $(G,\cdot,\circ)$, one can associate two Lie algebras:
@@
A post-Lie algebra $(\mathfrak{g},[\cdot,\cdot]_{\cdot},[\cdot,\cdot]_{\circ},\triangleright)$ is integrable if there exists an LSB $(G,\cdot,\circ)$ such that
@@
Remark 3.
Let $(\mathfrak{g},[\cdot,\cdot]_{\cdot},[\cdot,\cdot]_{\circ},\triangleright)$ be an integrable post–Lie algebra (PLA). Then $(\mathfrak{g},[\cdot,\cdot]_{\cdot},[\cdot,\cdot]_{\circ},\triangleright)$ is integrable by a connected LSB. Indeed, if $D(G,\cdot,\circ)=(\mathfrak{g},[\cdot,\cdot]_{\cdot},[\cdot,\cdot]_{\circ},\triangleright)$ for some (possibly disconnected) LSB $(G,\cdot,\circ)$, let $G_{0}$ be the identity component of $G$ (as a manifold). Since the two group structures share the same underlying manifold, $G_{0}$ is a Lie subgroup of both $(G,\cdot)$ and $(G,\circ)$, and the $\lambda$-maps preserve $G_{0}$. Hence $(G_{0},\cdot,\circ)$ is a connected LSB, and since $T_{e}G_{0}=T_{e}G$ we have
$D(G_{0},\cdot,\circ)=D(G,\cdot,\circ)=(\mathfrak{g},[\cdot,\cdot]_{\cdot},[\cdot,\cdot]_{\circ},\triangleright).$
@@
Consider the pre-Lie algebra $\left(\mathbf{M}_{n}(\mathbb{R}),[\cdot,\cdot]_{\cdot}=0,[\cdot,\cdot]_{\circ},\ \triangleright\right)$, where $X,Y]_{\circ}:=XY-YX$ is the usual matrix commutator, and the post-Lie product is defined by
@@
$x,y]_{\cdot}=0,\qquad[x,y]_{\circ}=xy-yx=0,$
so the underlying Lie algebras are abelian. The post-Lie product becomes $x\triangleright y=xy$, the standard multiplication in $\mathbb{R}$. Suppose, for contradiction, that this structure integrates to an LSB which is necessarily $(\mathbb{R},+,+)$. For such a trivial brace, we have
@@ 2.4. Affine actions
    There exists a Lie group structure $\cdot_{G}$ on $G$ such that $(G,\cdot_{G},\circ_{G})$ is an LSB;
@@
(1) $\Rightarrow$ (2): Suppose $(G,\cdot_{G},\circ_{G})$ is an LSB. Then the map
@@
where $\rho(g)=(x,f)$ and $e_{K}$ denotes the identity element of $(K,\cdot_{K})$. It follows that $\rho(G)$ is a regular subgroup of $\mathrm{Aff}(K,\cdot_{K})$. Since $\rho$ is an injective homomorphism of Lie groups it follows by [26, Prop. 7.17] and by Remark 1 that $\rho(G)$ is a closed subgroup of $\mathrm{Aff}(K,\cdot_{K})$.
Now, by Proposition 2.5, every regular Lie subgroup of $\mathrm{Aff}(K,\cdot_{K})$ corresponds to an LSB structure $(K,\cdot_{K},\circ_{K})$, and the group $\rho(G)$ is isomorphic to $(K,\circ_{K})$. Let $\Psi:\rho(G)\to(K,\circ_{K})$ be such an isomorphism. Then the composition
@@
which turns $G$ into a Lie group and makes $(G,\cdot_{G},\circ_{G})$ an LSB isomorphic to $(K,\cdot_{K},\circ_{K})$. The result follows. ∎
@@
    For every connected and simply-connected non-linear Lie group $(K,\cdot_{K})$, there exists a real connected linear Lie group $(G,\circ)$ and a simply transitive affine action $\rho:(G,\circ_{G})\longrightarrow\mathrm{Aff}(K,\cdot_{K})$.
@@
where $\mathrm{Aut}_{0}(G,\cdot)$ denotes the connected component of the identity in the group of Lie group automorphisms of $G$. Since $G$ is simply-connected, by [34, Thm. 3.18.13] it admits a Levi-Malcev decomposition $(G,\cdot)=R\rtimes S,$ where $R$ is the radical of $G$ and $S$ is a semisimple subgroup. Thus, in the terminology in [36], the subgroup $R$ is a nucleus, i.e. a closed connected and simply-connected normal, solvable Lie subgroup of $G$ such that $G/R$ is reductive. Since $R$ is invariant under the action of $\mathrm{Aut}_{0}(G)$, it follows by [36, Lemma 2] that there exists a faithful representation of $G$ which can be extended to a (possibly non-faithful) representation of the semidirect product $(G,\cdot)\rtimes\mathrm{Aut}_{0}(G,\cdot)$. Now consider the short exact sequence
$1\rightarrow(G,\cdot)\rightarrow(G,\cdot)\rtimes\mathrm{Aut}_{0}(G)\rightarrow\mathrm{Aut}_{0}(G)\rightarrow 1.$
Since $(G,\cdot)$ admits a faithful linear representation that extends to the semidirect product, and $\mathrm{Aut}_{0}(G,\cdot)$ is linear, being a Lie subgroup of $\mathrm{GL}(\mathfrak{g})$, where $\mathfrak{g}=\mathrm{Lie}(G,\cdot)$, we may apply [36, Lemma 1] to conclude that $\mathrm{Aff}_{0}(G,\cdot)$ is linear. ∎
Our second lemma shows how an LSB structure can be lifted to the universal covering of a connected LSB.
@@
[…]
@@
[…]
$\tilde{a}\circ(\tilde{b}\cdot\tilde{c})]\cdot[(\tilde{a}\circ\tilde{b})\cdot\tilde{a}^{-1}\cdot(\tilde{a}\circ\tilde{c})]^{-1}=\tilde{e}$ (7)
@@
i.e. $\pi\tilde{f}=\pi c_{\tilde{e}}=e$. Indeed, for $c_{\tilde{e}}$ it is obvious while for $\tilde{f}$:
@@
[…]
@@
[…]
@@
(S1) Let $(G,\cdot,\circ)$ be an LSB, with $(G,\cdot)$ connected simply-connected and linear. By Proposition 2.5, there exists a regular subgroup $H\leq\mathrm{Aff}(G,\cdot)$ such that $H\cong(G,\circ)$. Since $G$ is connected, it follows that $H\leq\mathrm{Aff}_{0}(G,\cdot)$. As $\mathrm{Aff}_{0}(G,\cdot)$ is linear (by Lemma 3.1), $H$ is also linear. Therefore, $(G,\circ)$ is linear as well.
@@
(R2) Let $(G,\cdot,\circ)$ be a connected Lie skew brace such that $(G,\circ)$ is semisimple and let $(\mathfrak{g},[\cdot,\cdot]_{\cdot},[\cdot,\cdot]_{\circ})$ be the corresponding PLA. Then the Lie algebra $(\mathfrak{g},[\cdot,\cdot]_{\circ})$ is semisimple as well. By adapting the proof of [12, Th. 3.3] to the real setting, [^2] Although [12, Th. 3.3] is stated over $\mathbb{C}$, the same rigidity argument can be transferred to the real case by passing to the complexifications and then descending to the real forms. we conclude that $(\mathfrak{g},[\cdot,\cdot]_{\cdot})\cong(\mathfrak{g},[\cdot,\cdot]_{\circ})$ as Lie algebras. Therefore, $(G,\cdot)$ and $(G,\circ)$ are locally isomorphic Lie groups. ∎
@@
and define $G:=H/N$, denoting the group operation on $G$ by $\circ$. We represent cosets in $G$ as $x,y,\theta]$, with $\theta\in\mathbb{S}^{1}$. As shown in [23, Section 4.8], the group $G$ is not linear. Now consider the following subgroups of $G$:
@@
It is readily verified that $G_{1}$ is a closed normal subgroup of $G$, $G_{2}$ is a closed subgroup, and $G_{1}\cap G_{2}=\{e\}$. Therefore, $G$ is the semidirect product $G=G_{1}\rtimes G_{2}$. Now endow the same underlying set $G=(\mathbb{R}\times\mathbb{S}^{1})\times\mathbb{R}$ with the direct product operation $\cdot$. According to Example 4, the pair of operations $(\cdot,\circ)$ defines a connected but non-simply-connected LSB, where the group $(G,\cdot)$ is linear, although $(G,\circ)$ is not. This shows that the assumption of simple-connectedness is crucial for the validity of point (S1).
@@
Because $G_{0}$ is connected, condition (S2) of Theorem 1.1 tells us that the group $(G_{0},\circ)$ is solvable. Next, $(G,\cdot)$ is compact and nilpotent, and $G_{0}$ is an open subgroup, so $G_{0}$ has finite index in $(G,\cdot)$. Therefore $|\overline{G}|<\infty$, and $(\overline{G},\cdot,\circ)$ is a finite left skew brace. Nilpotency is preserved by quotients, hence $(\overline{G},\cdot)$ is finite and nilpotent. By [31, Cor. 2.23] the multiplicative group $(\overline{G},\circ)$ is solvable. Since the class of solvable groups is closed under extensions, the group $(G,\circ)$ is also solvable. This completes the proof  [^3] Motivated by the discrete left skew brace setting (cf. [24]), one may call a subset $I\subseteq(G,\cdot,\circ)$ an ideal if it is a closed Lie subgroup satisfying (i) $I\trianglelefteq(G,\cdot)$, (ii) $I\trianglelefteq(G,\circ)$, (iii) $\lambda_{a}(I)=I$ for every $a\in G$. The argument above shows that $G_{0}$ is indeed such an ideal. The closedness of $I$ is precisely what guarantees that the quotient $G/G_{0}$ carries a natural manifold structure.. ∎
@@
(F1) Let $(G,\cdot)$ be a connected, simply connected, non-linear Lie group. Our objective is to endow $G$ with a second Lie group structure $\circ$ such that $(G,\cdot,\circ)$ forms an LSB, and the group $(G,\circ)$ is linear.
@@
In order to get (9) we use again the Levi–Malcev decomposition. So let $R\trianglelefteq G$ be a connected and simply connected solvable Lie subgroup, and $S\leq G$ be a connected and simply connected semisimple Lie group, such that
@@
Since $KAN=S\leq G$, we have $ANK=KAN$, so $r\in KAN$. But $R\cap KAN=\{e\}$, thus $r=e$, implying $uv=an\in AN$. However, $uv\in K$, and $K\cap AN=\{e\}$, hence $uv=e$, and so $x=e$. Therefore, $G_{1}\cap G_{2}=\{e\}$. Next, we prove that both $G_{1}$ and $G_{2}$ are linear. Note that $G_{1}=RUV=(RU)V.$ Since $R\trianglelefteq G$ and $R\cap U=\{e\}$, we have $RU\cong R\rtimes U$, which is connected, simply connected, and solvable. Also, $V$ is compact and therefore real linearly reductive. Since $UV$ normalizes $R$, it follows that $RU\trianglelefteq G_{1}$, and hence $G_{1}=(RU)\rtimes V$. By [21, Th. 16.2.7], such a group is linear. Moreover, $G_{2}=AN$ is connected, simply connected, and solvable, so it is linear by the same result.
(F2) Let $(G,\circ)$ be a connected and simply connected solvable Lie group. Combining Theorem 2.13 with Theorem 2.12, one can construct an LSB $(G,\cdot,\circ)$ such that the Lie group $(G,\cdot)$ is nilpotent.
@@
To justify each entry of the table in Theorem 1.1, we examine every pair $(G,\cdot)$ and $(G,\circ)$ in turn, constructing explicit non-trivial LSB structures for the boxes marked ✓ and proving non-existence or triviality for those marked with a dash (-).